OpenStudy (anonymous):

Each pail of plaster covers 90 square feet of ceiling. What is the least number of pails of plaster you would need to buy to cover the ceiling of a square room with walls 14 feet long?

OpenStudy (anonymous):

3 4 5 or 6

OpenStudy (anonymous):

first find the area of ceiling which is =14*14=196 square feet now 1 plaster cover=90 square feet so plaster required=196/90

OpenStudy (gorv):

196/90>2 so we require 3 plasters